Assume that the quantity theory of money holds and that velocity is constant at 5.0. Output...

  1. Assume that the quantity theory of money holds and that velocity is constant at 5.0. Output is fixed at its full-employment value of 10 000, and the price level is 2.0.

  1. Determine the real demand for money.

  1. The government fixes the nominal money supply at 5000. With output fixed at its fullemployment level and with the assumption that prices are flexible, what will be the new price level?

  1. What will be the price level if the government increases the nominal money supply to 6000?
Answer #1

Solution:-

A) PY=MV

2 * 10,000 = M * 5

20,000 = M * 5

M = 20,000 / 5

M = 4,000 (Nominal Money Supply)

Real money supply = Nominal money supply / Price level

                               = 4,000 / 2

                               = 2,000

B) When nominal money supply = 5,000

5,000 x 5 = P x 10,000

25,000 = 10,000P

P = 25,000/10,000 = 2.5

B) When nominal money supply = 6,000

6,000 x 5 = P x 10,000

30,000 = 10,000P

P =30,000/10,000 = 3